# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v4

@@ VIII Experiments
In this section, we present the comparative experimental results of FedFair and FedFDP with multiple baseline methods on three datasets: MNIST [22], FashionMNIST [48], and CIFAR10 [21]. We provide a detailed description of the experimental evaluation of FedFair and FedFDP in terms of performance, fairness, heterogeneity, and scalability. In particular, we conduct ablation studies on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$, and noise multiplier $\sigma$. In all tables, $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are better.
We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
VIII-A Default Settings
We compared our algorithms against six baseline methods. For FedFair, the baseline algorithms were left unmodified; for FedFDP, DP was added to each baseline algorithm to enable a fair comparison.
  1. 1.
    FedAvg[33]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. 2.
    SCAFFLOD[19]: SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. 3.
    FedProx[27]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. 4.
    FedDyn[2]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. 5.
    ALI-DPFL[31]: ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. 6.
    q-FFL[28]: q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings [30, 24] to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ [30, 46]. We use a 4-layer CNN architecture [33] which consists of two convolutional layers and two fully connected layers as model architecture.
Hyperparameters. For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
TABLE IV: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
Datasets MNIST FashionMNIST CIFAR10
FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
TABLE V: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
Herterogeneity Scalability
@@ VIII-A Default Settings
[…]
@@
VIII-B Result of FedFair
Performance and Fairness. Table II presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%. A plausible explanation is that, as the training process progresses, FedFair increasingly prioritizes fairness, resulting in a reduction in $\eta_{i}$. Compared to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods [55].
Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
Figure 4: Impact of different $C$ in FedFDP algorithm.
Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
Heterogeneity. Table III shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table III indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
VIII-C Results of FedFDP
Performance and Fairness. Table IV displays the experimental results of the FedFDP algorithm compared to baseline methods. Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
Heterogeneity. Table V shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table V demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline. Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
Impact of Different $\lambda$. We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm. As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
Impact of Different $C$. We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
@@ VIII-C Results of FedFDP
Different $\epsilon$. We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table VI. For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations. At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$. The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets. In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
Impact of Different $\sigma$. We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication. As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence. When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value. As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
IX Conclusion
@@
Appendix A Appendix
A-A Proof of Theorem 1
@@
$\mathbf{v}_{t+1}=\mathbf{w}_{t}\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (30)
A-A1 Key Lemma
@@
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 2, this implies:

### v6

@@ 7.3 Robust in Different Settings
Table 5: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
Heterogeneity Scalability
@@
Heterogeneity. Table 4 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively. Table 5 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table 4 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts. Table 5 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline. Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
Figure 2: Impact of different $\lambda$ in FedFDP algorithm.
Figure 3: Impact of different $C$ in FedFDP algorithm.
Figure 4: Impact of different $\sigma$ in FedFDP algorithm.
7.4 Performance under Different Hyper-parameters
In response to RQ3, we performed extensive hyper-parameter experiments. In Table 6, we observed results under fixed privacy budgets $\epsilon=\{1.0,2.0,3.0,4.0\}$. Table 7 presents experiments aiming at fixed target accuracies $\{80\%,85\%,90\%,95\%\}$. Fig. 2 explores the existence of an optimal $\lambda$, with results consistent with those in Section 5. The influence of the clipping norm $C$ is investigated in Fig. 3, and the effect of the noise multiplier $\sigma$ is examined in Fig. 4.
Impact of Different $\lambda$. In Section 5.1, we derive a closed-form solution for $\lambda^{*}$, which, however, relies on constants or upper bounds (e.g., $L$, $\mu$, $G$, $\Gamma$) whose specific values change with different models, datasets, and loss functions. Therefore, we conduct the experiment in Fig. 2 to investigate the practical range of $\lambda^{*}$. We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm. As depicted in Fig. 2, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
Impact of Different $C$. We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Fig. 3, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range {0.01, 0.1, 1, 5, 10}. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
Impact of Different $\sigma$. We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $\{1,1.5,2,2.5,3\}$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T\in\{6,115,268,463,708\}$ rounds of communication. As shown in Fig. 4, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence. When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value. As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
Table 6: The average accuracy ($\%$) and fairness ($\Psi$) on different privacy budget ($\epsilon$) setting at FashionMNIST dataset.
@@ 7.4 Performance under Different Hyper-parameters
Impact of Different $\epsilon$. We conducted experiments with $\epsilon\in\{1,2,3,4\}$ on the FashionMNIST dataset, as detailed in Table 6. For $\epsilon\in\{1,2,3\}$, we allowed larger values of $\epsilon$ to support more iterations. At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$. The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets. In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
Table 7: The privacy budget ($\epsilon$) and fairness ($\Psi$) on different target accuracy setting at MNIST $\mathrm{Dir}(0.05)$.
Acc. 80% 85% 90% 95%
$\epsilon(\downarrow)$ $\Psi(\downarrow)$ $\epsilon(\downarrow)$ $\Psi(\downarrow)$ $\epsilon(\downarrow)$ $\Psi(\downarrow)$ $\epsilon(\downarrow)$ $\Psi(\downarrow)$
FedAvg 2.43 2.10e9 3.02 8.90e9 3.45 1.30e10 4.89 1.30e11
SCAFFOLD 2.55 3.30e9 3.15 1.10e10 3.67 3.70e10 5.32 1.10e11
FedProx 3.02 1.30e9 2.98 1.30e10 4.31 5.80e10 4.89 2.50e11
FedDyn 2.68 2.20e9 3.55 7.80e9 3.98 1.10e10 5.15 3.30e11
ALI-DPFL 2.85 3.10e9 2.98 1.30e10 3.67 2.30e10 5.32 9.80e10
q-FFL 2.68 9.80e8 3.15 9.70e9 4.12 8.80e10 4.04 1.02e11
FedFDP 2.25 6.60e8 2.70 1.20e9 3.05 5.80e9 3.77 2.80e10
$\epsilon$ at Different Target Accuracy. TABLE 7 presents experiments on MNIST under the $\mathbf{Dir}(0.05)$ partition, reporting the privacy budget $\epsilon$ and fairness measure $\Psi$ at the point of achieving accuracy targets $\{80\%,85\%,90\%,95\%\}$. As the target accuracy increases (from $80\%$ to $95\%$), the privacy budget ($\epsilon$) for most methods rises slightly (e.g., FedAvg increases from 2.43 to 3.02), while fairness ($\Psi$) deteriorates significantly (e.g., FedAvg rises from 2.10×109 to 1.3×1011), highlighting the intensified trade-off between privacy and fairness under higher accuracy demands. FedFDP demonstrates exceptional performance in balancing these objectives: At the high $95\%$ accuracy target, its $\epsilon$ (2.70) is lower than all other methods, and its $\Psi$ (1.20×109) is significantly lower than competitors by an order of magnitude. At the $80\%$ accuracy target, FedFDP achieves the lowest $\Psi$ (6.60×108) and the lowest $\epsilon$ (2.25) in the entire table, validating its comprehensive superiority.
Figure 5: Tradeoff between privacy and fairness on $60\%$ accuracy.
Tradeoff between Privacy and Fairness. With the target accuracy fixed at $60\%$ on both MNIST and FashionMNIST, we investigate the trade-off between fairness and privacy. As shown in Fig. 5, as the privacy budget $\epsilon$ increases, the fairness metric $\Psi$ decreases significantly on both datasets, indicating a clear trade-off between privacy and fairness under the given accuracy constraint. The noise introduced by the privacy mechanism, while protecting data, amplifies performance disparities across groups, whereas relaxing privacy constraints creates room for fairness optimization. This monotonically decreasing relationship reveals the tripartite trade-off characteristics among privacy, fairness, and accuracy.
8 Limitations and Future Work
Despite the promising results achieved by FedFDP in balancing fairness, privacy, and utility, there are several limitations in the current study that open avenues for future research.
First, regarding heterogeneity, our current evaluation primarily focuses on statistical heterogeneity (Non-IID data) simulated via Dirichlet partitioning. We have not yet deeply explored model heterogeneity, where clients may possess different network architectures due to varying storage or memory capacities. Furthermore, the challenges associated with cross-domain data—where feature spaces or distributions differ fundamentally across clients—remain unaddressed. Future work could investigate adapting the fairness-aware gradient clipping strategy to support heterogeneous model architectures and cross-domain federated learning scenarios.
Second, the current framework assumes that all selected clients participate successfully in the training process (full participation). However, in practical real-world deployments, device heterogeneity (e.g., varying computational power) and unstable network conditions (e.g., bandwidth limits and high latency) often lead to the problem of stragglers. These stragglers can significantly delay the synchronization process in the global aggregation phase. In this work, we have not specifically optimized FedFDP for such system-level challenges. Future iterations of FedFDP could explore asynchronous update mechanisms or robust client selection protocols to mitigate the impact of stragglers and communication dropouts.
9 Conclusion
@@
Acknowledge
This work is supported by National Natural Science Foundation of China Key Program (Grant No. 62132005), Natural Science Foundation of Shanghai (Grant No. 22ZR1419100), and CAAI-Huawei MindSpore Open Fund (Grant No. CAAIXSJLJJ-2022-005A).
Appendix 0.A Appendix
In the beginning, Table 8 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
Table 8: Summary of main notations
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model at client $i$ at round $t$
$\mathbf{w}^{*}$ Optimal model that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
NN Total number of clients
$q$ batch sample ratio
$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equals to $|D_{i}|\cdot q$
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original clipping bound for loss
$||\cdot||$ $L_{2}$-norm
0.A.1 Convergence Analysis
Assumption 1
$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{\top}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 2
$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{\top}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 3
Theorem 0.A.1
where:
  • $A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}=\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}=\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}=\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive Equation (29). For the detailed proof, please refer to Appendix 0.A.2.
0.A.2 Proof of Theorem 0.A.1
@@ 0.A.1 Convergence Analysis
$\mathbf{v}_{t+1}=\mathbf{w}_{t}-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (31)
0.A.2.1 Key Lemma
@@
  • $A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}=\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}=\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}=\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Let $\Delta_{t}=\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 2, this implies: